OpenStudy (anonymous):

What is the value of the expression when a = 6 and b = 4?

OpenStudy (cutiecomittee123):

there isnt enough info to help you, what expression are we dealing with?

OpenStudy (anonymous):

14a-2(3b+4)

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

24

OpenStudy (anonymous):

whoops

OpenStudy (anonymous):

ok Thank you

OpenStudy (anonymous):

84-2(16)=84-32=52

OpenStudy (anonymous):

could u help me with a nother Plaese

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Which word phrases can be modeled by the expression? Choose exactly two answers that are correct. 8 − 3n A. eight decreased by the product 3 and a number B. the difference of 8 and the quotient of 3 and a number C. the difference of eight and the product of 3 and a number D. eight less than 3 times a number

OpenStudy (anonymous):

D m8 :)

OpenStudy (anonymous):

ya its d

OpenStudy (anonymous):

exactly 2

OpenStudy (anonymous):

there needs to be 2

OpenStudy (anonymous):

no one sec its a

OpenStudy (anonymous):

so d and a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which variable expression represents the word phrase? seven decreased by the sum of a number and four A. 7 – m + 4 B. 7 – (m + 4) C. (m + 4) – 7 D. 7 – (m – 4)

OpenStudy (anonymous):

you're giving her wrong answer buddy

OpenStudy (anonymous):

and this one to

OpenStudy (anonymous):

what

OpenStudy (anonymous):

eight less than 3 times a number is 3n-8 not 8-3n

OpenStudy (anonymous):

oh Whops sorry About that ya i just relized that